# Picocavity-enhanced Raman Spectroscopy of Physisorbed H2 and D2 Molecules

### v2

@@ Acknowledgements.
To compare with the above-mentioned studies, we performed TERS measurements while exposing D2 to a clean Ag(111) surfaces, as shown in Fig. S2. The gas pressure was set to Torr, comparable with the EELS study [11]. The ortho-D2-derived peaks appeared at 16 L and then saturated at 40 L, while para-D2 was not detected. The weak intensity at low exposures may be due to molecular diffusion along the surface at low coverages and the weak contribution of neighboring molecules to the TERS signal [see Fig. 2(c) in the main text]. It should be noted that even if the gas pressure and exposure values are comparable, the molecular coverage varies depending on the chamber configuration. Nevertheless, we confirmed that ortho-H2 and para-D2 were not detected at any pressures and exposures we used (0.8–2 Torr; 102–103 L). The HREELS observation of the metastable isomer at 6 K only at high exposures [11] may be attributed to be hydrogen multilayers, where the second layer does not interact with the Ag surface. For our LT-TERS measurement, in contrast, only a monolayer would be formed at the higher temperature, 10 K. Another possibility that the metastable isomer were not detected in TERS is that the LSPR in the gap is accelerating the para-to-ortho conversion of the molecule under the tip. Although far-field 532-nm irradiation did not affect the conversion rate [15], the strong electric field localized in the STM junction could interact with the nuclear spin of the adsorbate.
@@
To verify the contribution of LSPR to the TERS intensity, we examined the excitation wavelength () dependence. Figure S3(a) shows a STML spectrum recorded over a clean Ag(111) surface with an Ag tip. The LSPR results in a broad spectral feature in the visible range with a maximum at 560 nm. After recording the STML, we exposed the surface to the D2 gas, conducting TERS measurements with = 532 and 633 nm [Figs. S3(b) and S3(c), respectively]. The peak intensity of each mode significantly changes with , whereas the Raman shift value is unchanged. The vibrational intensity is weakened with = 633 nm, indicating the corresponding Raman scattering is not efficiently enhanced by the LSPR because the peak position ( = 779 nm) is located in the weak resonance range of the LSPR [Fig. S3(a)]. In contrast, both the rotational and vibrational peaks become intense with = 532 nm because they are more resonant with the LSPR.
@@
To obtain the waterfall plots shown in Fig. 3 of the main text, we conducted tip approach-and-retraction procedures while recording the TERS spectra [54, 41]. Figure S10(a) is the full dataset for the H2 “vib” mode shown in Fig. 3(b) of the main text. During recording TERS spectra, the tip height was lowered stepwise and then the tip was retracted from the surface. The symmetrical waterfall plot with respect to the closest tip height indicates the non-destructive measurement in the tip approach-and-retraction process. Figure S10(b) shows the horizontal line profiles of the waterfall plot for the H2 “vib” mode at the points indicated by the arrows in the same color. The three spectra indicates the red-shift and broadening of the H2 “vib” peak as the gap distance decreases.
@@
To determine the experimental peak position shift as a function of the tip height in Fig. 4(e) of the main text, the line profile of the waterfall plots at each tip height [Fig. 3 in the main text; see also Fig. S10(b)] was fitted by a Gaussian peak with a constant background. The top and bottom panels of Fig. S11 shows the peak position and width, respectively, as a function of the tip height for the rotational (a) and vibrational (b) modes. The error bars indicate the standard deviation of the fitting. As the tip approaches, the peak intensity increases, reducing the error. The point at which the error becomes below 2 cm-1 by the tip approach was used as the reference point [the arrows in the top panel of Fig. S11(a) and (b)] and the peak position at this reference point was set to the origin of the peak shift (0 cm-1 of the top panel). For the H2 and D2 “vib” modes [top panel of Fig. S11(b)], the plots after the reference points were used as the experimental plots shown in Fig. 4(e) of the main text. In Fig. 4(e) of the main text, the experimental tip height pm is aligned to the tip–sample distance nm for the calculations. Although the accurate determination of the experimental tip–sample gap distance is difficult, this alignment is reasonable; as described in Sect. I-E, pm is comparable with pm and nm.
As shown in Fig. S11(a), the frequency and width of the H2 and D2 “rot” peaks are almost unchanged by the tip. In contrast, the bottom panel Fig. S11(b) shows the strong peak broadening for the H2 “vib” mode [see also Fig. S10(b)] whereas the peak width for D2 “vib” is almost unchanged. A possible origin of the peak broadening is the anharmonicity, which can modify the lifetime of molecular vibrations [71]. It is also entirely possible that if neighboring molecules influence the experimental signal [see Fig. 2(c) in the main text], they would lead to the high-frequency shoulder of the vibration transition for H2, as these molecules would still lie on the surface [Fig. 3(c)] and would therefore present a higher than the ones on the tip, which instead provide the most enhanced signal; however, this effect may not be pronounced because the coverage of is less than that of D2 (see Sect. I-C). Another possible origin if the movement of the molecule in the junction; because the double-well potential barrier is very small [see Fig. 4(d) in the main text], the H2 molecule near the tip apex can move to a metastable site near the surface or adjacent adsorption sites on the surface [18] by a small voltage bias or plasmonic field applied in the STM junction. This movement instantaneously occurring during the acquisition time of the Raman spectrum can also modify the peak shape of the vibrational mode. We expect that the coexistence of these factors leads to the characteristic peak width depending on the tip height.

### v3

@@ Acknowledgements.
To compare with the above-mentioned studies, we performed TERS measurements while exposing D2 to a clean Ag(111) surfaces, as shown in Fig. S2. The gas pressure was set to Torr, comparable with the EELS study [11]. The ortho-D2-derived peaks appeared at 16 L and then saturated at 40 L, while para-D2 was not detected. The weak intensity at low exposures may be due to molecular diffusion along the surface at low coverages and the weak contribution of neighboring molecules to the TERS signal [see Fig. 2(c) in the main text]. It should be noted that even if the gas pressure and exposure values are comparable, the molecular coverage varies depending on the chamber configuration. Nevertheless, we confirmed that ortho-H2 and para-D2 were not detected at any pressures and exposures we used (0.8–2 Torr; 102–103 L). The HREELS observation of the metastable isomer at 6 K only at high exposures [11] may be attributed to be hydrogen multilayers, where the second layer does not interact with the Ag surface. For our LT-TERS measurement, in contrast, only a monolayer would be formed at the higher temperature, 10 K. Another possibility that the metastable isomer were not detected in TERS is that the localized surface plasmonic resonance (LSPR) in the gap is accelerating the para-to-ortho conversion of the molecule under the tip. Although far-field 532-nm irradiation did not affect the conversion rate [15], the strong electric field localized in the STM junction could interact with the nuclear spin of the adsorbate.
@@
To verify the contribution of LSPR to the TERS intensity, we examined the excitation wavelength () dependence. Figure S10(a) shows a STML spectrum recorded over a clean Ag(111) surface with an Ag tip. The LSPR results in a broad spectral feature in the visible range with a maximum at 560 nm. After recording the STML, we exposed the surface to the D2 gas, conducting TERS measurements with = 532 and 633 nm [Figs. S10(b) and S10(c), respectively]. The peak intensity of each mode significantly changes with , whereas the Raman shift value is unchanged. The vibrational intensity is weakened with = 633 nm, indicating the corresponding Raman scattering is not efficiently enhanced by the LSPR because the peak position (a wavelength of the Raman scattering of 779 nm) is located in the weak resonance range of the LSPR [Fig. S10(a)]. In contrast, both the rotational and vibrational peaks become intense with = 532 nm because they are more resonant with the LSPR.
@@
To obtain the waterfall plots shown in Fig. 3 of the main text, we conducted tip approach-and-retraction procedures while recording the TERS spectra [54, 41]. Figure S11(a) is the full dataset for the H2 “vib” mode shown in Fig. 3(b) of the main text. During recording TERS spectra, the tip height was lowered stepwise and then the tip was retracted from the surface. The symmetrical waterfall plot with respect to the closest tip height indicates the non-destructive measurement in the tip approach-and-retraction process. Figure S11(b) shows the horizontal line profiles of the waterfall plot for the H2 “vib” mode at the points indicated by the arrows in the same color. The three spectra indicates the redshift and broadening of the H2 “vib” peak as the gap distance decreases.
@@
To determine the experimental peak position shift as a function of the tip height in Fig. 4(e) of the main text, the line profile of the waterfall plots at each tip height [Fig. 3 in the main text; see also Fig. S11(b)] was fitted by a Gaussian peak with a constant background. The top and bottom panels of Fig. S12 shows the peak position and width, respectively, as a function of the tip height for the rotational (a) and vibrational (b) modes. To display the experimental peak-shift values , the reference point should be specified. Ideally is the furthest tip height at which the peak was detected. However, since using a too weak peak as the reference value causes large errors in the plot, we adopted as the tip height at which the error value below 2 cm-1 based on the peak fitting. The error bars in Fig. S12 indicate the standard deviation of the fitting. As the tip approaches, the peak intensity increases, reducing the error. As shown in the arrows in Fig. S12, we set , , , pm for H2 rot, D2 rot, H2 vib, and D2 vib, respectively, determining the origin of each peak shift. For the H2 and D2 “vib” modes [top panel of Fig. S12(b)], the plots at were used as the experimental curves shown in Fig. 4(e) of the main text. In Fig. 4(e) of the main text, the experimental tip height pm is aligned to the tip–sample distance nm for the calculations. Although the accurate determination of the experimental tip–sample gap distance is difficult, this alignment is reasonable; as described in Sec. I-E, pm is comparable with pm and nm.
As shown in Fig. S12(a), the frequency and width of the H2 and D2 “rot” peaks are almost unchanged by the tip. In contrast, the bottom panel Fig. S12(b) shows the strong peak broadening for the H2 “vib” mode [see also Fig. S11(b)] whereas the peak width for D2 “vib” is almost unchanged. A possible origin of the peak broadening is the anharmonicity, which can modify the lifetime of molecular vibrations [64]. It is also entirely possible that if neighboring molecules influence the experimental signal [see Fig. 2(c) in the main text], they would lead to the high-frequency shoulder of the vibration transition for H2, as these molecules would still lie on the surface [Fig. 3(c)] and would therefore present a higher than the ones on the tip, which instead provide the most enhanced signal; however, this effect may not be pronounced because the coverage of H2 is less than that of D2 (see Sec. I-C). Another possible origin if the movement of the molecule in the junction; because the double-well potential barrier is very small [see Fig. 4(d) in the main text], the H2 molecule near the tip apex can move to a metastable site near the surface or adjacent adsorption sites on the surface [18] by a small voltage bias or plasmonic field applied in the STM junction. This movement instantaneously occurring during the acquisition time of the Raman spectrum can also modify the peak shape of the vibrational mode. We expect that the coexistence of these factors leads to the characteristic peak width depending on the tip height.